OpenStudy (shamil98):

average rates f(x) = x^3 + 1 over the interval [2,3] so ((2)^3+1) - ((3)^3 +1) / 2 - 3 9 - 28 / -1 -19/-1 = 19 am i correct? ..

OpenStudy (anonymous):

Is this the f(b)-f(a)/b-a stuff?

OpenStudy (shamil98):

oh wait its f(x2) - f(x1)

OpenStudy (shamil98):

but the result is the same.

OpenStudy (shamil98):

28 - 9 / 1 = 19

OpenStudy (anonymous):

b-a though

OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

B-A

OpenStudy (anonymous):

b-a lets see if you figure it out

OpenStudy (shamil98):

brainfarted there, i read the formula wrong.

OpenStudy (anonymous):

<3

OpenStudy (shamil98):

((3^3) + 1) - ((2^3) + 1 ) / 3 -2 = 28 - 9 / 1 = 19 thanks lol. ^.^

OpenStudy (anonymous):

Yeah there we go

OpenStudy (anonymous):

You know I had no clue and guessed that formula..?

OpenStudy (shamil98):

good memory o.o this from limits lol

OpenStudy (shamil98):

so like calc 1 chapter 1 haha

OpenStudy (anonymous):

Ah right, been so long since I've seen it lol